At Guevarra clan reunion,4 1/2 kg of spaghetti was left,if there are 6 families,how much each family can take home equally?


1
Expert's answer
2020-10-29T20:45:21-0400

(4 1/2)/6 = 9/2


(9/2)/6 = (9/2)(1/6)


= 9/12


= 3/4 kg for each family